1 0
Add a comment Improve this question Transcribed image text
Answer #1

Using Ampere's law, for closed loop:

\(\oint B \cdot d s=\mu_{0} * I_{n e t}\)

\(\mu_{0}=\) Magnetic constant \(=4^{*} \mathrm{pi}^{*} 10^{\wedge}-7\)

\(I_{\text {net }}=\) Net current in the window frame \(=3.0 \mathrm{~A}\)

So,

\(\oint B . d s=\mu_{0} * I_{n e t}=4 \pi * 10^{-7} * 3.0=3.8 * 10^{-6} \mathrm{~T} . \mathrm{m}\)

Correct option is \(\mathrm{D}\).